In aqueous solution, hypobromite ion, BrO -, reacts to produce bromate ion, BrO 3-, and bromide ion, Br -, according to the following chemical equation.

3 BrO-(aq) → BrO3-(aq) + 2 Br-(aq)

A plot of 1/[BrO -] vs. time is linear and the slope is equal to 0.056 M -1s -1. If the initial concentration of BrO - is 1.0 M, how long will it take one-half of the BrO- ion to react?

12 s

18 s

5.6 × 10-2 s

8.9 s
Answer #1

Answer

18s

Explanation

1/[BrO-] vs time graph is straight line , so this is second order reaction

For second order reaction

half life , t1/2 = 1/(k[A]0)

where,

k = rate constant which equal to slope, 0.056M-1s-1

[A]0 = Initial concentration of reactant, 1.0M

substituting values

t1/2 =  1/(0.056M-1s-1 × 1.0M)

= 18 s

So,

time required to reduce the concentration of BrO- to one -half of the initial concentration = 18s